Assume that we want to construct a confidence interval. Do one of the following, as appropriate: (a) find the critical valueHow do you determine t or z? what should the choice be for this question?
Answer #1

Sol

Given

n = 287

Mean = 29.3

S.D = 7.5

As n is greater than or equal to 30

287 >= 30

Z table is used

Confidence interval = 90% = 0.90

Alpha = 1 - 0.90 = 0.1

Alpha/2 = 0.1 / 2 = 0.05

Z ( 0.05) from z table shows a value of 1.64

Therefore

Z alpha/2 = 1.64